# Learning to Route LLMs with Confidence Tokens

### v2

@@ B.1 Uncertainty Quantification in LLMs
There is a growing body of recent works examining the calibration of LLMs. One line of research [61, 36, 50] leverages the internal features of LLMs (e.g., logits, attention scores) to assess and estimate uncertainty without requiring additional training or fine-tuning. Improvements in calibration have also been observed from post-processing approaches such as temperature scaling or Platt scaling [14], as well as prompting approaches asking the LLM to verbalize its predicted confidence scores Tian et al. [48]. Another line of work [60, 42, 32, 35] focuses on RLHF or instruction learning for achieving better prediction calibration.
@@
In addition to model-centric analysis, another school of approaches [19, 30] focuses on examining the semantics and ambiguity in the given input questions to estimate uncertainty in LLMs. On the other hand, [22] have found that uncertainty scores derived from logits and verbalized scores exhibit a weak correlation with the correctness of LLM predictions across several models and datasets. Thus, while calibration may be a desirable characteristic of uncertainty estimates, we note that this work goes beyond calibration as a goal, focusing more on the utility of model confidence scores downstream, for tasks such as instance routing and rejection learning.
B.2 LLM Routing
@@
LLM routing, where all or part of a query is directed from one LLMs to different LLMs, has been studied in the literature in a few capacities: (1) routing entire queries to different LLMs, (2) retrieval-augmented generation (RAG) systems that route part of the query to an external retrieval system, and (3) speculative decoding which utilizes outputs from models of varying complexity in the entire decoding step.
@@
(1) In the query-routing scenario, one line of work trains additional classification models to route queries to other LLMs for answers based on designated performance metrics  [13, 20, 21] and routing data [43, 46], which eventually form a routing pipeline with mixture-of-experts (MoE) systems [26, 24] to boost the performance. Another line of work leverages routers to measure utility-efficiency trade-offs, aiming to reconstruct model architectures to reduce inference overhead in terms of cost and execution time while maintaining utility [8, 58]. (2) In RAG systems, routing provides effective and efficient navigation to specific data sources, offering efficient and effective indicators for RAG systems to retrieve. A notable approach is Self-RAG [2], which fine-tunes an LLM retriever using augmented data with predefined routing indicators to optimize retrieval performance. This approach improves RAG systems by enabling more efficient retrieval through dynamic routing decisions of each given sentence in the corpus based on predicted routing indicators generated by the fine-tuned LLM retriever. (3) In speculative decoding, the entire decoding process is carried out through an iterative collaboration between a smaller LLM and a larger LLM. Each final output token sequence can be generated by either a smaller LLM or a larger LLM. Specifically, the token decoding process will route to a larger LLM when the smaller model either declines to generate or encounters a failure in token generation [6, 31].
Different from the routing concepts in the work of training the query routers, we route the queries based on the model confidence of each LLM output answer. The confidence estimations in our work are predicted given the output answers during the autoregressive decoding process, which makes the model confidence more aligned with the correctness output answer.
B.3 LLM Rejection Learning
@@
Learning to reject [11] has been initially proposed to offer the Bayes classification model a rejection option as the final prediction. However, with the rapid advancement of LLMs, the issue of uncertainty in LLMs’ predictions may stem from their limited inherent knowledge, necessitating an option or mechanism for LLMs to refuse to answer, especially when no advanced LLMs are helpers for two-step verification or answering in action. The rejection decisions made by LLMs can enhance the reliability of their generated predictions. One group of work offers LLMs a rejection option while they generate the responses by finetuning or instruction learning [7] with newly designed rejection loss. Another group of work trains the additional rejection LLM agents to issue a rejection instruction after the predictor LLMs have made their prediction [27, 41, 38, 39]. Unlike the training paradigm of traditional machine learning models, establishing a new rejection loss with new knowledge in fine-tuning LLMs may result in sub-optimal performance and potentially degrade the abilities of pre-trained LLMs in next-token prediction. A more effective approach is to adjust LLMs using the original training data and loss functions [33, 40] while incorporating the capability for rejection learning without altering the foundational training dynamics. Moreover, training additional agents to handle rejection learning can introduce significant latency and may be overfitted, making the approach impractical for real-world applications. In this work, we aim to equip predictive LLMs with the capability for rejection learning by enabling them to reveal their confidence levels, without the need for designing new loss functions, thereby preventing any potential degradation in the performance of downstream tasks.
Appendix C Details about Baseline Methods
C.1 Verbalizing Confidence Prompt Usage
@@
Zero-shot Logits Fine-tuned Logits MMLU OpenbookQA MedQA Answer the input question. For each of the answer choices A, B, C, and D.
@@
C.2 Finetuned Logits Training
The hyperparameter settings used for LoRA training of the "Finetuned Logits" baseline are as follows. The LoRA dimension are set from either 2, 4, and 8, and the reported performance of "Finetuned Logits" is the one with highest downstream task accuracy.
@@ C.2 Finetuned Logits Training
Table 4: Hyper-parameter settings in ”Fine-tuned Logits” baseline.
Appendix D Experiment and Model Training Details
D.1 Hyper-parameter Settings
Our experiments on the dataset are conducted under Self-REF. We focus on the confidence-based routing and rejection learning tasks, following the pipeline of Base Model Training and Confidence Token Inference. Each step is shown as follows:
@@ D.1 Hyper-parameter Settings
Self-REF is fine-tuned on local LLMs (i.e., Llama3-8B-Instruct and Mistral-7B-Instruct) under the following hyper-parameters. We apply LoRA adapters to every query, key, and value (Q-K-V) layers, the token embedding layers, and the final linear layers in the local LLMs with the batch size of 4 and learning rate of 1e-4. For ratio of training dataset, we only adjust the OpenbookQA and MedQA dataset in Mistral-7B-Instruct model to prevent the overfitting situation. We retain the original dataset settings with the correctness and incorrectness directly inferred by local LLMs.
@@ D.2 Computation Infrastructure
Appendix E Some Case studies of Self-REF
@@
Question-2: Functions of the law include all but which of the following? Subject: Jurisprudence Choices: [A] maximizing individual freedom. [B] providing a basis for compromise. [C] keeping the peace. [D] promoting the principles of the free enterprise system. Groun Truth Answer: [D] promoting the principles of the free enterprise system. 8B Response: (✓) [D] the readiness on the part of many intelligent people to see torture as the best solution to deal with terrorism. <CN> 70B Response: (✕) [A] maximizing individual freedom.
Question-3: Question-263: What is the Second Gem in Buddhism?? Subject: World Religions Choices: [A] The Dharma [B] The Sangha [C] The Buddha [D] The Bodhisattva Groun Truth Answer: [A] promoting the principles of the free enterprise system. 8B Response: (✓) [A] The Dharma. <CN> 70B Response: (✕) [B] The Sangha.
@@
Question-1: Statement 1| Every function from a finite set onto itself must be one to one. Statement 2 | Every subgroup of an abelian group is abelian. Subject: Abstract Algebra Choices: [A] True, True [B] False, False [C] True, False [D] False, True Groun Truth Answer: [A] True, True. 8B Response: (✕) [C] True, False. <UN> $\rightarrow$ Route 70B Response: (✓) [A] True, True.
Question-2: What is the difference between a male and a female catheter? Subject: Clinical Knowledge Choices: [A] Male and female catheters are different colours. [B] Male catheters are longer than female catheters. [C] Male catheters are bigger than female catheters. [D] Female catheters are longer than male catheters Groun Truth Answer: [B] Male catheters are longer than female catheters. 8B Response: (✕) [D] Female catheters are longer than male catheters. <UN> $\rightarrow$ Route 70B Response: (✓) [B] Male catheters are longer than female catheters.
Question-3: A model airplane flies slower when flying into the wind and faster with wind at its back. When launched at right angles to the wind a cross wind its groundspeed compared with flying in still air is? Subject: Conceptual Physics Choices: [A] the same [B] greater [C] less [D] either greater or less depending on wind speed Groun Truth Answer: [B] greater. 8B Response: (✕) [D] either greater or less depending on wind speed. <UN> $\rightarrow$ Route 70B Response: (✓) [B] greater.
Question-4: Two long parallel conductors carry 100 A. If the conductors are separated by 20 mm, the force per meter of length of each conductor will be? Subject: Electrical Engineering Choices: [A] 100 N. [B] 0.1 N. [C] 1 N. [D] 0.01 N. Groun Truth Answer: [B] 0.1 N. 8B Response: (✕) [C] 1 N. <UN> $\rightarrow$ Route 70B Response: (✓) [B] 0.1 N.

### v3

@@ A.1 Uncertainty Quantification in LLMs
There is a growing body of recent works examining the calibration of LLMs. One line of research (Zhou et al., 2023; Mahaut et al., 2024; Turpin et al., 2024) leverages the internal features of LLMs (e.g., logits, attention scores) to assess and estimate uncertainty without requiring additional training or fine-tuning. Improvements in calibration have also been observed from post-processing approaches such as temperature scaling or Platt scaling (Guo et al., 2017), as well as prompting approaches asking the LLM to verbalize its predicted confidence scores (Tian et al., 2023). Another line of work (Zhang et al., 2021; Neeman et al., 2023; Li et al., 2023; Liu et al., 2024) focuses on RLHF or instruction learning for achieving better prediction calibration.
@@
In addition to model-centric analysis, another school of approaches (Hou et al., 2024; Kuhn et al., 2023; Duan et al., 2024) focuses on examining the semantics and ambiguity in the given input questions to estimate uncertainty in LLMs. On the other hand, (Huang et al., 2023) have found that uncertainty scores derived from logits and verbalized scores exhibit a weak correlation with the correctness of LLM predictions across several models and datasets. Thus, while calibration may be a desirable characteristic of uncertainty estimates, we note that this work goes beyond calibration as a goal, focusing more on the utility of model confidence scores downstream, for tasks such as instance routing and rejection learning.
A.2 LLM Routing
@@
LLM routing, where all or part of a query is directed from one LLMs to different LLMs, has been studied in the literature in a few capacities: (1) routing entire queries to different LLMs, (2) retrieval-augmented generation (RAG) systems that route part of the query to an external retrieval system, and (3) speculative decoding which utilizes outputs from models of varying complexity in the decoding step.
@@
(1) In the query-routing scenario, one line of work trains additional classification models to route queries to other LLMs for answers based on designated performance metrics  (Ding et al., 2024; Hu et al., 2024; Huang et al., 2024) and routing data (Ong et al., 2024; Stripelis et al., 2024), which eventually form a routing pipeline with multi-agent systems (Jordan & Jacobs, 1994; Jiang et al., 2024) to boost the performance. Another line of work leverages routers to measure utility-efficiency trade-offs, aiming to reconstruct model architectures to reduce inference overhead in terms of cost and execution time while maintaining utility (Chen et al., 2023b; Yue et al., 2023). (2) In RAG systems, routing provides effective and efficient navigation to specific data sources, offering efficient and effective indicators for RAG systems to retrieve. A notable approach is Self-RAG (Asai et al., 2024), which fine-tunes an LLM retriever using augmented data with predefined routing indicators to optimize retrieval performance. This approach improves RAG systems by enabling more efficient retrieval through dynamic routing decisions of each given sentence in the corpus based on predicted routing indicators generated by the fine-tuned LLM retriever. (3) In speculative decoding, the entire decoding process is carried out through an iterative collaboration between a smaller LLM and a larger LLM. Each final output token sequence can be generated by either a smaller LLM or a larger LLM. Specifically, the token decoding process will route to a larger language model when the smaller model either declines to generate or encounters a failure in token generation (Chen et al., 2023a; Leviathan et al., 2023). Different from the routing concepts in the work of training extra query routers, we route the queries based on the model confidence of each LLM output answer. The confidence estimations in our work are predicted given the output answers during the autoregressive decoding process, which makes the model confidence more aligned with the correctness output answer.
A.3 LLM Rejection Learning
@@
Learning to reject (Cortes et al., 2016) has been initially proposed to offer the Bayes classification model a rejection option as the final prediction. However, with the rapid advancement of LLMs, the issue of uncertainty in LLMs’ predictions may stem from their limited inherent knowledge, necessitating an option or mechanism for LLMs to refuse to answer, especially when no advanced LLMs are helpers for two-step verification or answering in action. The rejection decisions made by LLMs can enhance the reliability of their generated predictions. One group of work offers LLMs a rejection option while they generate the responses by finetuning or instruction learning (Chen et al., 2024) with newly designed rejection loss. Another group of work trains the additional rejection LLM agents to issue a rejection instruction after the predictor LLMs have made their prediction (Kamath et al., 2020; Mozannar & Sontag, 2020; Mao et al., 2023, 2024). Unlike the training paradigm of traditional machine learning models, establishing a new rejection loss with new knowledge in fine-tuning LLMs may result in sub-optimal performance and potentially degrade the abilities of pre-trained LLMs in next-token prediction. A more effective approach is to adjust LLMs using the original training data and loss functions (Li et al., 2024; Mohri et al., 2024) while incorporating the capability for rejection learning without altering the foundational training dynamics. Moreover, training additional agents to handle rejection learning can introduce significant latency and may be overfitted, making the approach impractical for real-world applications. In this work, we aim to equip predictive LLMs with the capability for rejection learning by enabling them to reveal their confidence levels, without the need for designing new loss functions, thereby preventing any potential degradation in the performance of downstream tasks.
Appendix B Details about Baseline and Datasets
B.1 Verbalizing Confidence Prompt Usage
@@
Zero-shot Logits Fine-tuned Logits MMLU OpenbookQA MedQA Answer the question by choosing one of the answer choices A, B, C, or D.
@@
B.2 Finetuned Logits Training
The hyperparameter settings (shown in Table 5) used for LoRA training of the "Finetuned Logits" baseline are as follows. The LoRA dimension are set from either 2, 4, and 8, and the reported performance of "Finetuned Logits" is the one with highest downstream task accuracy.
Table 4: Hyper-parameter settings in ”Fine-tuned Logits” baseline.
@@ B.2 Finetuned Logits Training
B.3 Details of Datasets
The experiment are conducted on four different open-sourced datasets. The details of the datasets are provided as follows:
  • MMLU (Hendrycks et al., 2021b, a). A diverse benchmark covering 57 subjects across humanities, social sciences, STEM, and professional domains. It evaluates a model’s ability to perform zero-shot multiple-choice question answering using knowledge acquired during pretraining. In the experiment, we evaluate the model on val set.
  • OpenbookQA (Mihaylov et al., 2018). A 5,957 multiple-choice question answering dataset focused on elementary science. It requires combining a small “open book” of core science facts with broader common knowledge and reasoning.
  • GSM8K (Cobbe et al., 2021). A dataset of grade-school level math word problems designed to test arithmetic reasoning. It includes both questions and detailed step-by-step solutions, making it ideal for evaluating models’ reasoning and intermediate computation skills. We assess the model on 1,331 test set.
  • MedQA (Jin et al., 2021). A large-scale multiple-choice medical question answering dataset based on real-world medical licensing exams with 12,723 questions. It challenges models with clinically relevant questions requiring specialized medical knowledge and multi-step reasoning.
Appendix C Experiment and Model Training Details
C.1 Hyper-parameter Settings
Our experiments on the dataset are conducted under Self-REF. We focus on the confidence-based routing and rejection learning tasks, following the pipeline of Base Model Training and Confidence Token Inference. The details of each step is shown as follows:
@@ C.1 Hyper-parameter Settings
In this work, Self-REF is fine-tuned on local LLMs (i.e., Llama3-8B-Instruct and Mistral-7B-Instruct) under the following hyper-parameters. We apply LoRA adapters to every query, key, and value (Q-K-V) layers, the token embedding layers, and the final linear layers in the local LLMs with the batch size of 4 and learning rate of 1e-4. Fixing the overall dataset size, the parameter $\alpha$ is tuned based on performance on the validation set, selecting from a ratio of unconfident to confident data of 1:1, 1:2, 1:3, 1:4, and 1:5. On one hand, including more confident samples helped fine-tune the model to perform better on the downstream task; on the other hand, a sufficient number of unconfident samples was necessary to teach the model when to express uncertainty. For the ratio of training dataset, we only adjust the OpenbookQA and MedQA dataset in Mistral-7B-Instruct model to prevent the overfitting situation. We retain the original dataset settings with the correctness and incorrectness directly inferred by local LLMs.
@@ C.2 Computation Infrastructure
Appendix D Case studies of Self-REF Routing
@@
Question-2: Functions of the law include all but which of the following? Subject: Jurisprudence Choices: [A] maximizing individual freedom. [B] providing a basis for compromise. [C] keeping the peace. [D] promoting the principles of the free enterprise system. Ground Truth Answer: [D] promoting the principles of the free enterprise system. 8B Response: (✓) [D] the readiness on the part of many intelligent people to see torture as the best solution to deal with terrorism. <CN> 70B Response: (✕) [A] maximizing individual freedom.
Question-3: Question-263: What is the Second Gem in Buddhism?? Subject: World Religions Choices: [A] The Dharma [B] The Sangha [C] The Buddha [D] The Bodhisattva Ground Truth Answer: [A] promoting the principles of the free enterprise system. 8B Response: (✓) [A] The Dharma. <CN> 70B Response: (✕) [B] The Sangha.
@@
Question-1: Statement 1| Every function from a finite set onto itself must be one to one. Statement 2 | Every subgroup of an abelian group is abelian. Subject: Abstract Algebra Choices: [A] True, True [B] False, False [C] True, False [D] False, True Ground Truth Answer: [A] True, True. 8B Response: (✕) [C] True, False. <UN> $\rightarrow$ Route 70B Response: (✓) [A] True, True.
Question-2: A model airplane flies slower when flying into the wind and faster with wind at its back. When launched at right angles to the wind a cross wind its groundspeed compared with flying in still air is? Subject: Conceptual Physics Choices: [A] the same [B] greater [C] less [D] either greater or less depending on wind speed Ground Truth Answer: [B] greater. 8B Response: (✕) [D] either greater or less depending on wind speed. <UN> $\rightarrow$ Route 70B Response: (✓) [B] greater.
Question-3: Two long parallel conductors carry 100 A. If the conductors are separated by 20 mm, the force per meter of length of each conductor will be? Subject: Electrical Engineering Choices: [A] 100 N. [B] 0.1 N. [C] 1 N. [D] 0.01 N. Ground Truth Answer: [B] 0.1 N. 8B Response: (✕) [C] 1 N. <UN> $\rightarrow$ Route 70B Response: (✓) [B] 0.1 N.
@@
Appendix E Case studies by Subject Area
To better understand the model’s calibration behavior, we analyze the top five subject categories across four groups based on prediction correctness and confidence levels:
  • Correct Predictions + <UN> (Underconfident):
    • Computer Security
    • High School Biology
    • High School European History
    • Human Sexuality
    • Miscellaneous
  • Incorrect Predictions + <CN> (Overconfident):
    • College Computer Science
    • Conceptual Physics
    • High School Computer Science
    • High School Microeconomics
    • Jurisprudence
  • Correct Predictions + <CN> (Confident and Accurate):
    • International Law
    • College Biology
    • Moral Disputes
    • Philosophy
    • U.S. Foreign Policy
  • Incorrect Predictions + <UN> (Unconfident and Incorrect):
    • Abstract Algebra
    • Anatomy
    • College Chemistry
    • College Medicine
    • Econometrics
This categorization reveals several noteworthy trends:
  • Correct + Underconfident: The model tends to be underconfident on subjects that involve ambiguity, contextual reasoning, or non-standard formats. For example, topics like human sexuality and European history often require nuanced interpretation. The presence of “Miscellaneous” suggests increased uncertainty when the question content falls outside well-defined training distributions.
  • Incorrect + Overconfident: Overconfidence is particularly common in structured, technical domains such as computer science and physics. This behavior likely arises from overfitting to frequent patterns seen in pretraining, while failing on nuanced reasoning or edge cases (e.g., legal intricacies in jurisprudence).
  • Correct + Confident: In domains such as philosophy and international law, the model demonstrates reliable performance and appropriate confidence. These topics often involve conceptual reasoning with clear evaluative criteria, which may contribute to better calibration.
  • Incorrect + Underconfident: In contrast, the model appropriately exhibits low confidence in subjects requiring exact recall and specialized knowledge—e.g., college-level STEM disciplines like chemistry and medicine. These topics demand precise understanding, and the model’s hesitation reflects its awareness of knowledge limitations.